# OpenGait: A Comprehensive Benchmark Study for Gait Recognition towards Better Practicality

### v1

@@ Abstract
Gait recognition, a rapidly advancing vision technology for person identification from a distance, has made significant strides in indoor settings. However, evidence suggests that existing methods often yield unsatisfactory results when applied to newly released real-world gait datasets. Furthermore, conclusions drawn from indoor gait datasets may not easily generalize to outdoor ones. Therefore, the primary goal of this work is to present a comprehensive benchmark study aimed at improving practicality rather than solely focusing on enhancing performance. To this end, we first develop OpenGait, a flexible and efficient gait recognition platform. Using OpenGait as a foundation, we conduct in-depth ablation experiments to revisit recent developments in gait recognition. Surprisingly, we detect some imperfect parts of certain prior methods thereby resulting in several critical yet undiscovered insights. Inspired by these findings, we develop three structurally simple yet empirically powerful and practically robust baseline models, i.e., DeepGaitV2, SkeletonGait, and SkeletonGait++, respectively representing the appearance-based, model-based, and multi-modal methodology for gait pattern description. Beyond achieving SoTA performances, more importantly, our careful exploration sheds new light on the modeling experience of deep gait models, the representational capacity of typical gait modalities, and so on. We hope this work can inspire further research and application of gait recognition towards better practicality. The code is available at https://github.com/ShiqiYu/OpenGait.
@@
OpenGait, Gait Recognition, Benchmark Study.
@@ I Introduction
Gait recognition has garnered increasing interest within the vision research community. It leverages physiological and behavioral characteristics observed in walking videos to authenticate individuals’ identities [1]. Compared with other biometrics, such as the face, fingerprint, and iris, gait patterns can be captured from a distance in uncontrolled settings, without necessitating physical contact. Additionally, as a natural walking behavior, gait is inherently difficult to disguise, rendering it theoretically robust against common subject-related variables, such as clothing, carrying items, and posture variations. These advantages position gait recognition as a promising solution for security applications, including identity authentication and suspect tracking [2].
@@
In previous studies, gait methods were primarily developed using private code repositories and heavily reliant on in-the-lab gait datasets, particularly CASIA-B [10] and OU-MVLP [11]. However, these datasets are partially limited, with CASIA-B comprising data from 124 subjects collected nearly 20 years ago, and OU-MVLP involving only cross-view changes. To accelerate real-world applications, this work considers building a unified evaluation platform, that covers various SoTA gait methods and emerging gait datasets captured in realistic environments, as highly desired. In line with this vision, we introduce a flexible and efficient gait recognition platform termed OpenGait. Sec. III-A will provide the highlight features of OpenGait. Thanks to its extensibility and reusability, OpenGait has been extended to new potentially influential repositories, such as Gait3D-Benchmark [9] [^1] https://github.com/Gait3D/Gait3D-Benchmark, FastPoseGait [12] [^2] https://github.com/BNU-IVC/FastPoseGait, and All-in-One-Gait [^3] https://github.com/jdyjjj/All-in-One-Gait. Moreover, OpenGait has also been widely adopted in two major international gait recognition competitions, i.e., HID [13], and GREW [8]. Notably, all top-10 winning teams at HID2022 [13] and 2023 [14] have utilized OpenGait as the codebase to develop new solutions.
@@
Using OpenGait as the foundation, we reproduce several SoTA methods, with some results presented in Fig. 1 (b). Furthermore, we reassess some commonly accepted conclusions by re-implementing the ablation study on recently curated outdoor gait datasets. Surprisingly, we find that the MGP branch proposed by GaitSet [7], the FConv proposed by GaitPart [15], the local feature extraction branch proposed by GaitGL [16], and the SMPL branch proposed by SMPLGait [9], do not demonstrate superiority on real-world gait datasets. With exhaustive analysis, we reveal several critical yet overlooked limitations of existing gait research, including insufficient ablation study, lack of outdoor evaluation, and the absence of a strong backbone, etc.
@@
Inspired by the above discoveries, we develop three structurally simple, experimentally powerful, and empirically robust baseline models for real-world gait recognition. Specifically, we introduce DeepGaitV2444We regard GaitBase, proposed by the conference version of this work [17], as an initial exploration of deep ResNet [18] for gait modeling and denote it as DeepGaitV1. Consequently, its subsequent iteration is labeled as DeepGaitV2., SkeletonGait, and SkeletonGait++, each utilizing the silhouette image, skeletal coordinates, and the combination of these two as input, accordingly standing for the appearance-based, model-based, and multi-modal methodology for gait pattern description. As illustrated in Fig. 1 (c), the silhouette and skeleton present two of the most popular gait inputs [^5] We count 78 gait papers released over the past 5 years, across top computer vision conferences such as CVPR, ICCV, ECCV, BMVC, WACV, ACCV, and related top conferences like AAAI, ACM MM, ICASSP, ICIP, and IJCB. underscoring their adoption. In terms of network design, this benchmark study prioritizes the use of widely accepted blocks and gait frameworks to enhance the practicality and applicability of our findings. Through extensive experiments, several critical insights regarding the modeling capacity of shallow vs. deep gait models and the representational ability of gait silhouette vs. skeleton has been provided. By addressing these key challenges straightly, our pragmatic solutions yield significant advancements as depicted in Fig. 1 (b). More details will be provided in Sec. V.
@@
In summary, this benchmark study contributes to gait research in three significant ways: a) OpenGait, a unified and extensible platform, is developed to facilitate the systematic analysis of gait recognition. b) We conduct a comprehensive experimental review of recent advancements in gait recognition, yielding insightful findings. c) We introduce three robust baseline models, while maintaining practicality and commonality, demonstrating significant performance enhancement.
@@
This work stems from our CVPR2023 highlight paper, OpenGait [17], and makes three significant advancements: a) We upgrade the GaitBase to DeepGaitV2 by thoroughly exploring the capability of shallow vs. deep models for gait modeling. b) We combine findings from our AAAI2024 publication [19], specifically SkeletonGait and SkeletonGait++, with DeepGaitV2, resulting in a more comprehensive baseline reconstruction study. c) We broaden the experimental scope to include recently released popular gait datasets, such as CCPG [20] and SUSTech1K [21]. By aiming to enhance the practicality of gait recognition, this paper offers a more systematic analysis compared to its conference counterpart.
@@ II Related Work
This section starts with a literature review, involving the aspects of the gait dataset, input modality, and popular method, providing a snapshot of recent developments in the field. Following this, we offer a brief review of other related works.
@@
Environment Dataset Train Set Test Set Cameras Variations Modalities Venue
@@
Constrained CASIA-B [10] 74 8,140 50 5,500 11 #CV, #CL, #BG Sil., RGB ICPR’06
OU-MVLP [11] 5,153 144,284 5,154 144,412 14 #CV Sil., Ske. IPSJ’18
CCPG [20] 100 8,187 100 8,095 10 #CV, #CL, #BG Sil., RGB CVPR’23
SUSTech1K [21] 200 5,988 850 19,228 12 #CV, #CL, #BG, etc Sil., RGB, Lidar CVPR’23
In-the-wild GREW [8] 20,000 102,887 6,000 24,000 882 Real-world Sil., Ske., Point ICCV’21
Gait3D [9] 3,000 18,940 1,000 6,369 39 Real-world Sil., Ske., Mesh CVPR’22
#ID and #Seq present the number of identities and sequences. #CV, #CL, and #BG are for camera viewpoints, clothing changes, and bag carrying.
Sil. and Ske. respectively refer to the silhouette and skeleton.
@@ II-A Gait Recognition Datasets
The large-scale data collection presents an essential premise for contemporary gait recognition research. As exhibited in Table I, several popular gait datasets can be roughly categorized into two groups: the constrained and in-the-wild gait datasets, where the former ones often require subjects to repeatedly walk along fixed routes with simulating real scenarios by introducing dressing and carrying changes, while the latter ones are captured from real-world scenarios naturally covering a wide range of real-world covariates. Even though the constrained gait datasets have imported more and more human-crafted complexities, formulating gait recognition through in-the-wild settings remains a relatively practical and challenging choice according to mainstream experimental results [22]. Before the release of this paper’s conference version [17], most existing works only verify their effectiveness on in-the-lab datasets, e.g., CASIA-B [10] and OU-MVLP [11], thus posing a high risk of vulnerability for practical usage. In this benchmark study, we evaluate several representative SoTA methods on emerging real-world gait datasets thus uncovering overlooked issues. In addition, the effectiveness of our three kinds of baseline models is verified on both constrained and in-the-wild gait datasets, convincingly demonstrating the generality and significance of our findings for gait recognition.
@@ II-B Typical Gait Modalities
As shown in Fig. 1 (c), the mainstream gait modalities, such as the binary silhouette, skeleton coordinates, RGB image, human mesh, and body parsing image, are primarily derived from cameras and represented by various data formats. During this process, diverse pretreatment operations [7, 23, 24, 21, 25] and end-to-end learning manners [26, 27] have been utilized to mitigate the influence of gait-irrelevant noises like color, texture, and background cues. Out of this trend, some studies propose novel gait modalities by incorporating emerging sensors such as LiDAR [21] and event cameras [28]. However, these sensors are currently less commonly found in CCTVs, making them temporarily unsuitable for large-scale surveillance applications. Even so, OpenGait supports all of the above gait modalities regardless of the sensor type.
In addition, Fig. 1 (c) further reveals that the binary silhouette and skeleton coordinates act as two of the most prevailing gait modalities in the latest literature. They both explicitly present the human body structural characteristics, e.g., the length, ratio, and movement of human limbs. Silhouettes, especially, have a more discriminative capacity by explicitly presenting the body shape information. When the gait recognition challenge evolves from constrained evaluations to in-the-wild applications, the gait models based on these two modalities have suffered different degrees of performance degradations [17, 29]. Except for recognition accuracy renaissances, this work further explores the cooperativeness and complementarity natures of these two gait modalities for ‘comprehensive’ gait pattern descriptions, thus proposing the multi-modal SkeletonGait++.
@@ II-C Gait Recognition Methods
For summary convenience, we roughly group gait recognition methods into three categories, i.e., the model-based, appearance-based, and multi-modal methods.
@@
Model-based Gait Recognition methods [23, 30, 27] tend to take the estimated underlying structure of the body as input, such as 2D/3D pose and SMPL [31] model. With extremely excluding visual clues, these gait modalities, typically parameterized as coordinates of body joints or customized vectors, are ideally ‘clean’ against factors like carrying and dressing items. In recent literature, PoseGait [23] used 3D body pose and hand-crafted structural features to overcome the changes in clothing. GaitGraph [30] introduced the graph convolutional network for 2D skeleton-based gait description. GPGait [29] developed a human-oriented transformation and a series of human-oriented descriptors to generate the unified pose representation with multiple features. Generally, the skeleton-based methods often struggle with non-shape-information situations and, as a result, usually perform unsatisfactorily on a variety of gait datasets. On the other hand, several SMPL-based methods have achieved significant advances on indoor OU-MVLP [11], e.g., Li et al. [27] fine-tuned a pre-trained human mesh recovery network to construct the end-to-end SMPL-based model and Xu et al[32] proposed an occlusion-aware human mesh model-based method to alleviate the partial occlusion problem within practical gait recognition. However, the competitiveness of this line of work on real-world gait datasets [8, 9] has not yet been verified exactly.
@@
Appearance-based Gait Recognition methods mostly learn gait features from binary silhouettes or RGB images and benefit from informative shape characteristics. With the boom of deep learning, most current appearance-based works focus on spatial feature extraction and temporal modeling. Specifically, GaitSet [7] innovatively regarded the gait sequence as a set and utilized a maximum function to compress frame-level spatial features. Thanks to its simplicity and effectiveness, GaitSet has become one of the most influential gait methods in recent years. GaitPart [15] carefully explored the local details of the input silhouette and modeled the temporal dependencies [15]. GaitGL [16] argued that the spatially global gait representations often neglect the details, and the local region-based descriptors cannot capture the relations among neighboring parts [16]. 3DLocal [33] wanted to extract limb features through 3D local operations at adaptive scales. DyGait [34] proposed to establish spatial-temporal representations of dynamic body parts. In a statistical sense, the appearance-based methods present the focus of the recent research.
@@
Multi-modal Gait Recognition methods typically incorporate multiple gait modalities as input, reflecting an emerging trend in gait pattern description. In recent literature, SMPLGait [9] exploited 3D information from the SMPL model to enhance the learning of gait appearance features. BiFusion [35] integrated body skeletons and silhouettes to capture the rich spatio-temporal features. ParsingGait [25] regarded the human parsing image as a form of fine-grained segmentation to improve the representational ability of the silhouette branch.
@@
In addition to the above, many successful model-agnostic gait frameworks promote gait research as well, e.g., GaitEdge [26] designed an edge-trainable silhouette to build the end-to-end gait recognition framework, GaitGCI [36] introduced the generative counterfactual intervention to alleviate the over-fitting problem within gait recognition, GaitSSB [37] collected millions of unlabelled gait sequences to pre-train the gait model by contrastive learning, LidarGait [21] built the first large-scale in-the-wild lidar-based gait method to explore precise 3D gait features from point clouds, and UDA [38] developed a novel fine-grained unsupervised domain adaptation method for robust gait recognition.
To act as a comprehensive and extensible gait recognition platform, the proposed OpenGait is compatible with almost all of the aforementioned methods.
Figure 2: The design principles of proposed codebase OpenGait. Seg is for segmentation, Rec is for recognition, and Br is for branch.

### v2

@@ Abstract
Gait recognition, a rapidly advancing vision technology for person identification from a distance, has made significant strides in indoor settings. However, evidence suggests that existing methods often yield unsatisfactory results when applied to newly released real-world gait datasets. Furthermore, conclusions drawn from indoor gait datasets may not easily generalize to outdoor ones. Therefore, the primary goal of this paper is to present a comprehensive benchmark study aimed at improving practicality rather than solely focusing on enhancing performance. To this end, we developed OpenGait, a flexible and efficient gait recognition platform. Using OpenGait, we conducted in-depth ablation experiments to revisit recent developments in gait recognition. Surprisingly, we detected some imperfect parts of some prior methods and thereby uncovered several critical yet previously neglected insights. These findings led us to develop three structurally simple yet empirically powerful and practically robust baseline models: DeepGaitV2, SkeletonGait, and SkeletonGait++, which represent the appearance-based, model-based, and multi-modal methodologies for gait pattern description, respectively. In addition to achieving state-of-the-art performance, our careful exploration provides new perspectives on the modeling experience of deep gait models and the representational capacity of typical gait modalities. In the end, we discuss the key trends and challenges in current gait recognition, aiming to inspire further advancements towards better practicality. The code is available at https://github.com/ShiqiYu/OpenGait.
@@
OpenGait; Gait recognition; Benchmarking; Biometrics.
@@ I Introduction
Gait recognition has garnered increasing interest within the vision research community. It leverages physiological and behavioral characteristics observed in walking videos to authenticate individuals’ identities [1]. Compared with other biometric features, such as the face, fingerprint, and iris, gait patterns can be captured from a distance in uncontrolled settings, without necessitating physical contact. Additionally, as a walking behavior, gait is inherently difficult to disguise, rendering it theoretically robust against common subject-related variables, such as clothing, carrying items, and posture variations. These advantages position gait recognition as a promising solution for security applications, including identity authentication and suspect tracking [2].
@@
In previous studies, the source code was typically shared in the authors’ repositories, and the methods relied heavily on in-the-lab gait datasets, particularly CASIA-B [10] and OU-MVLP [11]. However, these datasets have significant limitations compared to in-the-wild datasets. For instance, CASIA-B includes data from only 124 subjects collected nearly 20 years ago, while OU-MVLP involves only cross-view changes. To accelerate real-world applications, there is a pressing need to build a unified evaluation platform that encompasses various gait methods and emerging gait datasets captured in realistic environments. To this end, we introduce OpenGait, a flexible and efficient gait recognition platform, with its highlight features detailed in Sec. III-A. Thanks to its extensibility and reusability, OpenGait has been extended to some new influential repositories, such as Gait3D-Benchmark [9] [^1] https://github.com/Gait3D/Gait3D-Benchmark, FastPoseGait [12] [^2] https://github.com/BNU-IVC/FastPoseGait, and All-in-One-Gait [^3] https://github.com/jdyjjj/All-in-One-Gait. Moreover, OpenGait has also been widely adopted in two major international gait recognition competitions, i.e., HID [13], and GREW [8]. Notably, all top-10 winning teams at HID2022 [13] and 2023 [14] chose OpenGait as the codebase to develop their solutions.
@@
With the help from OpenGait, we efficiently reproduce several SoTA methods, with some results presented in Fig. 1 (b). Furthermore, we reassess some commonly accepted conclusions by re-implementing the ablation study on recently curated outdoor gait datasets. Surprisingly, we find that the MGP branch proposed by GaitSet [7], the FConv proposed by GaitPart [15], the local feature extraction branch proposed by GaitGL [16], and the SMPL branch proposed by SMPLGait [9], do not demonstrate superiority on real-world gait datasets. With exhaustive analysis, we reveal several critical yet overlooked limitations of existing gait research, including insufficient ablation study, lack of outdoor evaluation, and the absence of a strong backbone, etc.
@@
Inspired by the above discoveries, we develop three structurally simple, experimentally powerful, and empirically robust baseline models for real-world gait recognition. Specifically, we introduce DeepGaitV2444We regard GaitBase, proposed by the conference version of this work [17], as an initial exploration of deep ResNet [18] for gait modeling and denote it as DeepGaitV1. Consequently, its subsequent iteration is labeled as DeepGaitV2., SkeletonGait, and SkeletonGait++, each utilizing the silhouette image, skeletal coordinates, and the combination of these two as input, accordingly standing for the appearance-based, model-based, and multi-modal methodology for gait pattern description. As illustrated in Fig. 1 (c), the silhouette and skeleton present two of the most popular gait inputs [^5] We count 78 gait papers released over the past 5 years, across top computer vision conferences such as CVPR, ICCV, ECCV, BMVC, WACV, ACCV, and related top conferences like AAAI, ACM MM, ICASSP, ICIP, and IJCB. underscoring their adoption. In terms of network design, this benchmark study prioritizes the use of widely accepted modules and gait frameworks to enhance the practicality and applicability of our findings. Through extensive experiments, several critical insights regarding the modeling capacity of shallow vs. deep gait models and the representational ability of gait silhouette vs. skeleton has been provided. By addressing above challenges straightly, our pragmatic solutions yield significant advancements as depicted in Fig. 1 (b). More details will be provided in Sec. V.
Before the conclusion, this study presents a concise discussion on key trends and challenges in gait recognition under real-world scenarios. This section covers critical topics such as ideal gait input, handling noisy data, in-the-wild challenges, gait metric learning, and gait temporal modeling. By offering valuable insights into the current state of gait recognition and outlining a clear roadmap for future advancements, it aims to drive research toward better practicality.
@@
In summary, this benchmark study makes four key contributions: a) We develop OpenGait, a unified and extensible platform designed to facilitate the systematic analysis of gait recognition. b) We conduct a comprehensive experimental review of representative gait recognition methods, yielding insightful findings. c) We introduce three robust baseline models, while maintaining practicality and commonality, demonstrating significant performance enhancements. d) We provide a brief discussion section highlighting key trends and challenges to inspire future gait recognition research.
@@
This work stems from our CVPR 2023 highlight paper, OpenGait [17], and makes four significant advancements: a) We upgrade the GaitBase to DeepGaitV2 by thoroughly exploring the capability of shallow vs. deep models for gait modeling. b) We combine findings from our AAAI 2024 publication [19], specifically SkeletonGait and SkeletonGait++, with DeepGaitV2, resulting in a more comprehensive baseline reconstruction study. c) We broaden the experimental scope to include recently released popular gait datasets, such as CCPG [20] and SUSTech1K [21]. (d) We add a brief section discussing the key trends and challenges in gait recognition, incorporating feedback from the OpenGait community. Based on these significant changes, this paper offers a more systematic analysis aimed at improving the practicality of gait recognition compared to its conference counterpart.
@@ II Related Work
This section starts with a literature review that covers the gait datasets, input modalities, and popular methods. It aims to provide a snapshot of recent developments in the field.
@@
Environment Dataset Train Set Test Set Cameras Variations Modalities Year
@@
Constrained CASIA-B [10] 74 8,140 50 5,500 11 #CV, #CL, #BG Sil., RGB 2006
OU-MVLP [11] 5,153 144,284 5,154 144,412 14 #CV Sil., Ske. 2018
FVG [22] 136 - 90 - 1 #CV, #CL, etc RGB 2019
CCPG [20] 100 8,187 100 8,095 10 #CV, #CL, #BG Sil., RGB 2023
SUSTech1K [21] 200 5,988 850 19,228 12 #CV, #CL, #BG, etc Sil., RGB, Lidar 2023
CCGR [23] 600 908322 370 672295 33 #CV, #CL, #BG, etc Sil., RGB, Parsing, etc 2024
In-the-wild GREW [8] 20,000 102,887 6,000 24,000 882 Real-world Sil., Ske., Flow 2021
Gait3D [9] 3,000 18,940 1,000 6,369 39 Real-world Sil., Ske., Mesh 2022
Note: #ID and #Seq present the number of identities and sequences. #CV, #CL, and #BG are for camera viewpoints, clothing changes, and bag carrying. Sil. and Ske. respectively refer to the silhouette and skeleton.
@@ II-A Gait Recognition Datasets
The large-scale data collection presents an essential premise for gait recognition research. As exhibited in Table I, several popular gait datasets can be roughly categorized into two groups: the constrained and the in-the-wild ones. The former group often requires subjects to repeatedly walk along fixed routes with simulating real scenarios by introducing dressing and carrying changes. The latter ones should be captured from real-world scenarios and cover a wide range of real-world covariates. Even though the constrained gait datasets can include many human-crafted complexities, in-the-wild data is still essential to gait recognition [24]. Before the release of this paper’s conference version [17], many existing works only verify their effectiveness on constrained in-the-lab datasets, e.g., CASIA-B [10] and OU-MVLP [11], thus posing a high risk of vulnerability for practical usage. In this benchmark study, we evaluate several representative methods on emerging real-world gait datasets thus uncovering overlooked issues. In addition, the effectiveness of our three baseline models has been verified on both constrained datasets and in-the-wild ones. The results have convincingly demonstrated the generality and significance of our findings for gait recognition.
Beyond identification purposes, gait datasets for medical applications have gained increasing attention in recent literature. Notable examples include Scoliosis1K [25] for scoliosis classification, as well as PsyMo [26] and D-Gait [27] for psychological assessment. OpenGait has integrated support for some of these datasets and recognizes vision-based gait analysis as a promising direction for medical research.
@@ II-B Typical Gait Modalities
As shown in Fig. 1 (c), the mainstream gait modalities, such as the binary silhouette, skeleton coordinates, RGB image, human mesh, and body parsing image, are primarily derived from cameras and represented by various data formats. During this process, diverse pretreatment operations [7, 28, 29, 21, 30] and end-to-end learning manners [31, 32] have been utilized to mitigate the influence of gait-irrelevant noises like color, texture, and background cues. Out of this trend, some studies propose novel gait modalities by incorporating emerging sensors such as LiDAR [21] and event cameras [33]. However, these sensors have not been widely deployed in surveillance scenarios, making them unsuitable for large-scale surveillance applications currently. To ensure comprehensiveness, OpenGait supports all of the aforementioned gait modalities, regardless of sensor types.
In addition, Fig. 1 (c) further reveals that binary silhouettes and skeleton coordinates act as two of the most prevailing gait modalities in the latest literature. They both explicitly present the human body structural characteristics, e.g., the length, ratio, and movement of human limbs. For the rich body shape information, the silhouette has a better discriminative capacity in gait description. However, regardless of the modality employed, gait recognition methods will suffer accuracy degradation when tested in outdoor scenarios. [17, 34]. Beyond performance, we also try to explore the cooperativeness and complementarity natures of these two gait modalities by proposing a multi-modality method, SkeletonGait++.
@@ II-C Gait Recognition Methods
For convenience, we roughly group gait recognition methods into three categories, i.e., the model-based, appearance-based, and multi-modal methods.
@@
Model-based Gait Recognition methods [28, 35, 32, 36] tend to take the estimated underlying structure of the body as input, such as 2D/3D pose and SMPL [37] model. With extremely excluding visual clues, these gait modalities, typically parameterized as coordinates of body joints or customized vectors, are ideally clean against factors like carrying and dressing items. In the recent literature, PoseGait [28] employs 3D body pose and hand-crafted structural features to overcome the changes in clothing. GaitGraph [35] introduces the graph convolutional network for 2D skeleton-based gait description. In GPGait [34] a human-oriented transformation and a series of human-oriented descriptors are developed to generate the unified pose representation with multiple features. GaitPT [38] adopts a hierarchical transformer design, effectively extracting spatial and temporal movement features in an anatomically consistent manner. Generally, the skeleton-based methods often struggle with the absence of body shape and, as a result, usually perform unsatisfactorily on most gait datasets. On the other hand, several SMPL-based methods have achieved significant advances on indoor OU-MVLP [11], e.g., Li et al. [32] fine-tuned a pre-trained human mesh recovery network to construct the end-to-end SMPL-based model and Xu et al[39] proposed an occlusion-aware human mesh based method to alleviate the partial occlusion problem within practical gait recognition. However, model-based methods on real-world gait datasets [8, 9] have not yet been verified exactly.
@@
Appearance-based Gait Recognition methods mostly learn gait features from binary silhouettes or RGB images and can benefit from informative shape characteristics. With the boom of deep learning, most current appearance-based works focus on spatial feature extraction and temporal modeling. Specifically, in GaitSet [7] a gait sequence is innovatively regarded as a set and the max pooling is utilized to compress frame-level spatial features. Thanks to its simplicity and effectiveness, GaitSet has become one of the most influential gait methods in recent years. In GaitPart [15] the local details of the input silhouette are carefully explored and the temporal dependencies are modeled. In GaitGL [16] it is argued that the spatially global gait representations often neglect the details, and the local region-based descriptors cannot capture the relations among neighboring parts. 3DLocal [40] tries to extract limb features through 3D local operations at adaptive scales, and DyGait [41] tries to establish spatial-temporal representations of dynamic body parts. Generally speaking, appearance-based methods contribute the majority in the recent gait recognition.
@@
Multi-modal Gait Recognition methods typically incorporate multiple gait modalities as input, reflecting an emerging trend in gait pattern description. In recent literature, SMPLGait [9] exploits 3D information from the SMPL model to enhance the learning of gait appearance features. BiFusion [42] integrates body skeletons and silhouettes to extract the rich spatio-temporal features. ParsingGait [30] regards a human parsing image as a form of fine-grained segmentation to improve the representational ability of its silhouette branch.
Figure 2: The three main modules of OpenGait. In the Modeling Module, Seg is for segmentation, Rec is for recognition, and Br is for branch.
@@
In addition to the above, many successful model-agnostic gait frameworks promote gait research as well. For example, GaitEdge [31] employs edge-trainable silhouettes to build an end-to-end gait recognition framework, GaitGCI [43] introduces a generative counterfactual intervention to alleviate the over-fitting problem within gait recognition, GaitSSB [44] builds a pre-trained model with millions of unlabelled gait sequences by contrastive learning, LidarGait [21] collects the first large-scale in-the-wild lidar-based gait dataset to explore gait feature extraction from precise 3D point clouds, and UDA [45] develops a novel fine-grained unsupervised domain adaptation method for robust gait recognition.
To be noting that almost all of the aforementioned methods and some other recent methods have been implemented in OpenGait. OpenGait can help researchers to evaluate and compare different methods efficiently.